OpenStudy (asib1214):

A sample of a gas occupies 0.785L when the pressure is 243.4KPa and the temperature is 15C. Calculate the volume the gas will occupy at SATP.

OpenStudy (asib1214):

the formula i'm using is P1V1/T1 = P2V2/T2

OpenStudy (asib1214):

but i'm confused with V2 P@ and T2..............................

OpenStudy (asib1214):

@tester97

OpenStudy (tester97):

Oh i forgot this sorry honey

OpenStudy (asib1214):

its alright, stay cute!!!!

OpenStudy (tester97):

Thanks :)

OpenStudy (abmon98):

1litres-->0.001m^3 0.785litres-->x x=0.785*0.001=0.000785m^3 Use Ideal Gas Equation: PV=nRT R is a gas constant and is equivalent to 8.31 Temperature is changed from 15 degrees to kelvin by adding to 273 0.000785*243.3*10^3=n*8.31*(273+15) n=0.079 moles The standard temperature is 273 K (0° Celsius) and the standard pressure is 1 atm pressure. At STP, one mole of gas occupies 22.4 L of volume (molar volume). 22.4*number of moles=Volume 22.4*0.079=V

OpenStudy (asib1214):

this is how i did it. V1=0.785L P1=234.1KPa T1=15C P2=100KPa T2=25C P1V1/T1=P2V2/T2 234.1KPa X 0.785L/15C = 100KPa X V2/25C 25 cancels out from the right side and shifts to the left side because i'm trying to isolate the V2 25X183.76/25 = 100 KPa X V2 now 100 KPa cancels out from the right side and shifts to the left side 306.26/100KPa = V2 3.062L = V2 Check if i'm right guys???????

OpenStudy (asib1214):

i'm working with SATP and not STP

OpenStudy (asib1214):

@Somy

OpenStudy (anonymous):

gay-lusac and other stuff right?

OpenStudy (asib1214):

if you are referring to the gas laws then yes

OpenStudy (anonymous):

asib you are correct

OpenStudy (asib1214):

aight thanks for the feedback!!!!